Robinson Crusoe produces upper-loop product of $1000. He pays $750 in wages, $125 in interest, and $75 in rent. What must his profi t be? If three-fourths of Crusoe’s output is consumed and the rest invested, calculate Crusoeland’s GDP with both the product and the income approaches and show that they must agree exactly. 


1
Expert's answer
2021-11-22T09:55:39-0500

Profit is given by total revenue less all expenses(total cost) incurred


Robinson crusoe revenue is "\$ 1000"


Robinson crusoe expenses;


Rent "\$75"


Interest "\$125"


Wages "\$750"


Total expenses = "\$950"


Profit = 1000 - 950


Profit = "\$50"


GDP;


Product approach


GDP = C + I


Consumption is three quarters of 1000 = 750


Investment = 1000- consumption


Investment= 1000- 750


Investment = 250


Therefore GDP = 750 + 250


GDP= "\$1000"




INCOME APPROACH


GDP = Wages + Rent + Interest + Profit


GDP = 750 + 75 + 125 + 50


GDP = "\$1000"




Conclusion, both product approach and income gives the same GDP of $1000.